As a member of the Finance Department of Ranch Manufacturing Co., your supervisor has asked you to compute the appropriate discount rate to use when evaluating the purchase of new packaging equipment for the plant. Under the assumption that the firm’s present capital structure reflects the appropriate mix of capital sources for the firm, you have determined the market value of the firm’s capital structure as follows:
Source of Capital   Market values
Bonds. $4,000,000
Preferred stock. $2,000,000
Common stock.   $6,000,000

To finance the purchase, Ranch Manufacturing Co. will sell 10-year bond with a 7% yield to maturity. Preferred stock paying a $2 dividend can be sold for $25. Common stock for Ranch Manufacturing is currently selling for $55 per share & the firm paid a $3 dividend last year. Dividends are expected to continue growing at a rate of 5% per year into the indefinite future. If the firm’s tax ratio is 30%, what discount rate should you use to evaluate the equipment purchase?
Answer #1
1] The first step is to find the component costs:
After cost of debt = YTM*(1-t) = 7%*(1-30%) = 4.90%
Cost of preferred stock = Dividend/Price = 2/25 = 8.00%
Cost of common stock using the constant dividend growth model = D0*(1+g)/P0+g, where D0 = last dividend, g = growth rate in dividends and P0 = current price.
= 3*1.05/55+0.05 = 10.73%
2] WACC = Cost of debt*Weight of debt+Cost of preferred stock*Weight of preferred stock+Cost of equity*Weight of equity
The WACC is calculated in the following table:
Component Market value Weight Comonent cost WACC
Bonds $       40,00,000 33.33% 4.90% 1.63%
Preferred stock $       20,00,000 16.67% 8.00% 1.33%
Common stock $       60,00,000 50.00% 10.73% 5.36%
Total $   1,20,00,000 100.00% 8.33%
WACC = 8.33%
The discount rate to be used for evaluating the equipment purchase is the WACC of 8.33%.